OpenStudy (anonymous):

Solve the eqution below for the given variable. (x+3)^2/5=4... two answers

OpenStudy (anonymous):

(x+3)^2/5=4 so (x+3)^2 = 20 take sqrt of both sides x+3 = plus or minus sqrt20 and x = plus or minus sqrt20 - 3